# Automatically Analyzing Performance Issues in Android Apps: How Far Are We?

### v1

@@ Abstract.
Performance plays a critical role in ensuring the smooth operation of any mobile application, directly influencing user engagement and retention. Android applications are no exception. However, unlike functionality issues, performance issues are more challenging to discover as their root causes are sophisticated and typically emerge under specific payloads. To tackle this problem, researchers have dedicated substantial efforts to proposing automatic approaches for understanding, detecting, and resolving performance issues. Despite these endeavors, it still remains unknown what the status quo of Android performance analysis is, and whether existing approaches can indeed accurately reflect real performance issues. To fill this research gap, we conducted a systematic literature review followed by an explanatory study to explore relevant studies and real-world challenges. Our findings reveal that current tools have limited capabilities, covering only 17.50% of the performance issues. Additionally, existing datasets encompass only 27.50% of the issues and are very limited in size. We also show real-world issue patterns, underscoring the huge gap between the identified techniques and practical concerns. Furthermore, possible solutions are provided to guide future research towards achieving effective performance issue detection and resolution.
@@ 1. Introduction
With the increasing popularity of smartphones, the Android system has become one of the largest mobile platforms in the world, holding a 70.69% market share worldwide (iPhoneAndroidStatistics, ). The exponential growth of Android usage has seen an unprecedented expansion within the mobile app industry, with millions competing for users’ attention. In such a fiercely competitive circumstance, developers are under pressure to create new, fancy apps in the quickest manner. However, this emphasis on rapid development can sometimes come at the expense of app quality, particularly in terms of performance.
Unlike functionality issues that directly cause app malfunctions or failures, performance issues occur when the app functions correctly but its efficiency is low, resource usage is unreasonable, or user interaction is poor, significantly reducing user satisfaction and software reliability (Li2019DetectingAD, ; Liu2014CharacterizingAD, ; Gao2017EveryPC, ; Liu2018NavyDroidAE, ). These issues include, but are not limited to, excessive energy consumption (Banerjee2018EnergyPatchRR, ; Pathak2012WhatIK, ; Carette2017InvestigatingTE, ), resource leaks (Wu2016LightWeightIA, ; Guo2013CharacterizingAD, ; Lu2022DetectingRU, ; Yan2013SystematicTF, ), and responsiveness delays (Farooq2018RuntimeDroidRR, ; Kang2016DiagDroidAP, ; Lin2024AgingOG, ). For example, Liu et al. (Liu2014CharacterizingAD, ) demonstrate that wasteful energy consumption can occur in background threads that continue to update invisible GUIs or perform complex network computations, rapidly depleting battery life and degrading app performance. Similarly, Yan et al. (Yan2013SystematicTF, ) reveal resource leak defects associated with events such as screen rotation, app switching, pressing the back button, and file or database access.

### v2

@@ Abstract.
Performance issues in Android applications significantly undermine users’ experience, engagement, and retention, which is a long-lasting research topic in academia. Unlike functionality issues, performance issues are more difficult to diagnose and resolve due to their complex root causes, which often emerge only under specific conditions or payloads. Although many efforts haven attempt to mitigate the impact of performance issues by developing methods to automatically identify and resolve them, it remains unclear if this objective has been fulfilled, and the existing approaches indeed targeted on the most critical performance issues encountered in real-world settings. To this end, we conducted a large-scale comparative study of Android performance issues in real-world applications and literature. Specifically, we started by investigating real-world performance issues, their underlying root causes (i.e., contributing factors), and common code patterns. We then took an additional step to empirically summarize existing approaches and datasets through a literature review, assessing how well academic research reflects the real-world challenges faced by developers and users. Our comparison results show a substantial divergence exists in the primary performance concerns of researchers, developers, and users. Among all the identified factors, 57.14% have not been examined in academic research, while a substantial 76.39% remain unaddressed by existing tools, and 66.67% lack corresponding datasets. This stark contrast underscores a substantial gap in our understanding and management of performance issues. Consequently, it is crucial for our community to intensify efforts to bridge these gaps and achieve comprehensive detection and resolution of performance issues.
Android, Software Maintenance, Performance Issues
@@ 1. Introduction
With the increasing popularity of smartphones, the Android system has become one of the largest mobile platforms in the world, holding a 70.69% market share worldwide (iPhoneAndroidStatistics, ). The exponential growth of Android usage has seen an unprecedented expansion within the mobile app industry, with millions competing for users’ attention. In such a fiercely competitive circumstance, developers face pressure to create new, fancy apps in the quickest manner (Nikzad2014APEAA, ). However, this emphasis on speed often leads to a focus on functional requirements, while non-functional requirements, particularly performance, are overlooked—even though they are crucial to the overall quality of the apps (Hort2021ASO, ).
According to the classic SE textbook from Ian Sommerville, performance issues arise when software meets functional requirements but operates inefficiently or wastes resources, which can negatively affect its reliability (sommerville2011software, ). These issues are particularly critical in Android mobile devices, where resources like battery life, CPU, memory, and storage are limited (Guo2013CharacterizingAD, ; Liu2014CharacterizingAD, ; Rua2023ALE, ; Lee2019ImprovingEE, ). In Android apps, performance issues affecting efficiency, resource utilization, or app smoothness can drastically reduce user satisfaction and software reliability (Li2019DetectingAD, ; Liu2014CharacterizingAD, ; Gao2017EveryPC, ; Liu2018NavyDroidAE, ). These issues include, but are not limited to, excessive energy consumption (Liu2014CharacterizingAD, ; Banerjee2018EnergyPatchRR, ; Pathak2012WhatIK, ; Carette2017InvestigatingTE, ), redundant memory consumption (Guo2013CharacterizingAD, ; Song2021IMGDroidDI, ; Habchi2019TheRO, ), and responsiveness delays (Farooq2018RuntimeDroidRR, ; Kang2016DiagDroidAP, ; Lin2024AgingOG, ).